OpenStudy (annjewell1998):

You drove at 55 mph for four and a half hours. Which of the following equations will tell you how far you drove? Please help

OpenStudy (annjewell1998):

a D = (55)(4.5) b 55 = R (4.5) c 4.5 = 55 T d 4.5 = R (55)

OpenStudy (anonymous):

hint: distance = rate * time

OpenStudy (annjewell1998):

so D= 55*4 1/5?

OpenStudy (annjewell1998):

no its 55= R (4.5)